OpenStudy (anonymous):

What is the degrees Fahrenheit if the temperature is -60°C? Show all work. Remember to label your answer. If necessary, round your final answer to the nearest tenth. I already know that the answer is -76F but im not sure how i should explain my answer.

jimthompson5910 (jim_thompson5910):

F = (9/5)C + 32 ... start with the celsius to fahrenheit formula F = (9/5)(-60) + 32 ... plug in C = -60, then evaluate/simplify F = (9/5)(-60/1) + 32 F = (9*(-60))/(5*1) + 32 F = (-540)/(5) + 32 F = -108 + 32 F = -76

jimthompson5910 (jim_thompson5910):

As for the explanations, you're simply following PEMDAS (the order of operations) after you replace C with -60.

OpenStudy (anonymous):

thanks love.

jimthompson5910 (jim_thompson5910):

yw